OpenStudy (anonymous):

Evaluate the function y=cos(sin^-1(x)) geometrically to find the answer in radical form. Set your answer equal to y.

OpenStudy (anonymous):



OpenStudy (anonymous):

What is really important here is the equation: \[ \sin^2(x)+\cos^2(x)=1 \]

OpenStudy (anonymous):

thats true, i havent studied the trigonometric identities much yet.

OpenStudy (anonymous):

We get \[ \sin^2(x) = 1-\cos^2(x) \]

OpenStudy (anonymous):

\[\sin(x)\sin(x)=1-\cos(x)\cos(x)\]

OpenStudy (anonymous):

We also get \[ \cos^2(x) = 1-\sin^2(x) \]Taking square root of both sides: \[ \cos(x) = \sqrt{1-\sin^2(x)} \] \[ y=\cos(\sin^{-1}(x)) = \sqrt{1-\sin^2(\sin^{-1}(x))} \]Hmmmm.

OpenStudy (anonymous):

See what I did there?

OpenStudy (anonymous):

so \[y=\cos(\sin ^{-1}(x))\]

OpenStudy (anonymous):

eeeeeh, basically what you said. i dun wanna retype it .w. but yes, i think so.

OpenStudy (anonymous):

\[ \cos(\color{red}x) = \sqrt{1-\sin^2(\color{red}x)} \\ \cos(\color{red}{\sin^{-1}(x)}) = \sqrt{1-\sin^2(\color{red}{\sin^{-1}(x)})} \]

OpenStudy (anonymous):

I hope you are not color blind.

OpenStudy (anonymous):

and the whole thins is under the square root because in the original equation the inverse sine of x was the "x"...im not color blind :3